Calculate the molality of the a. 0.050 g CO2 in 652 g of H2O, b. 56 g NH3 in 200 g of H2O, c. 3.21 g C6H12 in 2.3 g of benzene (C6H6).
1
Expert's answer
2021-02-15T02:48:07-0500

a) Moles of "CO_2=" "0.050CO_2\times" "1 molCO_2\over 44.01gCO_2" "=0.0011molCO_2"

"1g=0.001L"

Liter of solution="0.652L"

Molarity"=" "Moles of solute\over liter of solution"

"=" "0.0011mol \over 0.652L" "0.00174mol\/L"


b) Moles of "NH_3=56gNH_3\times" "1mol NH_3\over 17.03gNH_3" "=3.289mol NH_3"

Liter of solution"=0.2L"

Molarity"=Moles of solute\over liter of solution"

"=3.289mol\over 0.2L" "=16.44mol\/L"


c) Moles of "C_6H_{12}=" "3.21gC_6H_{12}\times" "1molC_6H_{12}\over 84.16gC_6H_{12}" "=0.0381molC_6H_{12}"

Liter of solution"=0.0023L"

Molarity"=" "Moles of dilute\over liter of solution"

"=" "0.0381mol\over 0.0023L" "16.56 mol\/L"